# Exploring the statistical properties of the neutron-deficient 109In isotope with the Oslo method

### v1

@@ II.1 Experimental details and the Oslo method
Figure 1: (a) Experimental $p-\gamma$ coincidence data, (b) unfolded, and (c) primary matrices for 109In. Yellow dashed lines mark the neutron separation energy. The ranges of excitation and $\gamma$-ray energies used for the analysis with the Oslo method are shown with solid blue lines in (c). The gates on the ground state with spin-parity $9/2^{+}$ (D1) and a cluster of states with spin-parities $3/2^{-}$, $11/2^{+}$, $5/2^{+}$, $1/2^{+}$ (D2) used for the shape method are shown with red dashed lines in (c). Excitation-energy and $\gamma$-ray-energy bins are 124-keV wide.
@@
The extracted primary $\gamma$-ray distributions $P(E_{\gamma},E_{i})$ are further decomposed in terms of the $\gamma$ transmission coefficient $\mathcal{T}_{i\rightarrow f}$ (proportional to the GSF) and the NLD $\rho_{f}$ as
@@
As outlined in Ref. [39], the decomposition provides the functional forms of the solutions for the NLD and the $\gamma$ transmission coefficient defined in the general form as:
@@
With one set of solutions found, $\rho(E_{i}-E_{\gamma})$ and $\mathcal{T}(E_{\gamma})$, an infinite set of equivalent solutions $\tilde{\rho}(E_{i}-E_{\gamma})$ and $\tilde{\mathcal{T}}(E_{\gamma})$ can be constructed by applying the scaling factors $A$ and $B$ and modifying the slope of the resulting functions with the parameter $\alpha$. To single out the unique physical solutions for the NLD and the $\gamma$ transmission coefficient, these parameters must be constrained using additional experimental data.
II.2 Normalization of the nuclear level density and $\gamma$-ray strength function
The standard normalization procedure employed in the Oslo method to determine parameters $A$, $B$, and $\alpha$ makes use of the low-lying discrete levels, available in the compilation [48], and $s$- or $p$-wave neutron resonance data listed in the Reference Input Parameter Library [49] and the Atlas of Neutron Resonances [50]. This part of the analysis is described in great detail in earlier works, e. g., in [25, 28, 21]. In this section, we briefly summarize the key points of the normalization procedure.
@@ II.2 Normalization of the nuclear level density and $\gamma$-ray strength function
These procedures allow us to determine the scaling $A$ and the slope $\alpha$ of the NLD, where the latter also gives the slope of the GSF. As a final step, the scaling $B$ of the GSF is extracted using the average total radiative width $\langle\Gamma_{\gamma}\rangle$ from $s$- and/or $p$-wave neutron capture experiments [50, 49]. The expression for the average total radiative width $\langle\Gamma_{\gamma}(E_{x},J,\pi)\rangle$ from Ref. [62] [Eq. (3.1)] can be adapted to account for the states accessed through the capture of $\ell=0$ or $\ell=1$ neutrons. For example, for the $s$-wave case and a target nucleus with $J_{t}\neq 0$, the GSF $f(E_{\gamma})$ can be expressed through the value of $\langle\Gamma_{\gamma}\rangle$ as [25]:
$\langle\Gamma_{\gamma}\rangle=$ $\langle\Gamma(S_{n},J_{t}\pm 1/2,\pi_{t})\rangle=\frac{1}{2\rho(S_{n},J_{t}\pm 1/2,\pi_{t})}\times$
$\times\int_{E_{\gamma}=0}^{S_{n}}dE_{\gamma}E_{\gamma}^{3}f(E_{\gamma})\rho(S_{n}-E_{\gamma})\times$
@@
Here, we used the proportionality of the $\gamma$ transmission coefficient $\mathcal{T}(E_{\gamma})$ and the GSF $f(E_{\gamma})$, $\mathcal{T}(E_{\gamma})=2\pi E_{\gamma}^{3}f(E_{\gamma})$, for dipole transitions expected to dominate in the studied energy region [62]. Within the Oslo method, it is not possible to distinguish between magnetic ($M$) and electric ($E$) types of transitions, unless external experimental information on the $M1$ distribution is available. The GSF extracted with Eq. (9) is therefore the sum of $E1$ and $M1$ dipole strengths. The GSF is extrapolated to low (and in some cases high) $E_{\gamma}$ energies to calculate the integral in Eq. (9) as discussed in Ref. [25].
@@
II.3 Normalization of the 109In data
With neutron resonance data at hand, the normalization of both the NLD and the GSF becomes a relatively straightforward process. However, they are not accessible for many stable and unstable nuclei near the edges of the valley of stability, due to the lack of suitable target nuclei for neutron-capture experiments. In some cases, systematics for neighboring nuclei can be used to constrain the values of $\rho(S_{n})$ and $\langle\Gamma_{\gamma}\rangle$ [22, 21]. This and other assumptions made during the normalization process commonly result in comparatively large systematic uncertainties for both the NLD and the GSF.
The situation becomes even more challenging for odd-$Z$ nuclei with very few stable isotopes available; in such cases, the question is to what extent systematics from even-$Z$ neighboring nuclei can be reliably applied. This challenge was, for example, encountered in a recent analysis of 127Sb [28], which relied largely on neutron resonance data from the Sn and Te isotopic chains. Likewise, very little information is available on In isotopes (except for the stable 113In and the long-lived 115In), that could potentially be used in the normalization procedure for the neutron-deficient 109In. For the In isotopes, located one proton below the $Z=50$ closure, a solution similar to that adopted for 127Sb is suggested by the abundant neutron resonance data available for the Sn and Cd nuclei. The $\langle\Gamma_{\gamma}\rangle$ value for the odd-even 109In can be extracted in a similar way as described in Ref. [28], from the trend observed in experimental s-wave $\langle\Gamma_{\gamma}\rangle$ values for the even-odd Sn and Cd isotopes from Ref. [50]. All values, as well as the results of a $\chi^{2}$-fit with a quadratic function, are shown in Fig. 2.
@@ II.3 Normalization of the 109In data
To reduce the uncertainty in the slope shared by these functions, the so-called shape method was proposed in Ref. [63]. This technique has been successfully applied in different mass regions [63, 58, 45, 64] to reconstruct the general shape of the GSF (and, therefore, the slope of the NLD) from data on the direct decay to low-lying states. The number of counts $N$ associated with these decays is confined within a diagonal ($D$) in the primary matrix. The number of counts for a given $E_{i}$ bin can be written in a general form as [63]:
@@
where $f(E_{\gamma})E_{\gamma}^{3}$ is proportional to the $\gamma$-ray transmission coefficient, $\sigma(E_{i},J_{i})$ is the population cross-section for a state with spin $J_{i}$ at excitation energy $E_{i}$, and $g(E_{i},J_{i})$ is the intrinsic spin distribution given by Eq. (5). In cases when a diagonal contains more than one state, the first sum runs over all spins $J_{f}$ of the final states. As in the Oslo method, parity equipartition is adopted in Eq. (12), and only dipole transitions are taken into account. This restricts the applicability of the shape method to excitation energies where dipole transitions are dominant. As a next step, the population cross section is assumed to be independent of spin $J_{i}$ in the chosen region of excitation energies $E_{i}$. For dipole decays from a set of initial states to the final states with spins $J_{f}$, confined within the diagonal $D$ at the excitation-energy bin $E_{i}$ in the primary matrix, the GSF can therefore be expressed as:
@@
The functional form of the NLD extracted from the decomposition in Eq. (1) exhibits a clear BSFG trend, suggesting that this function (Eq. (3) in Ref. [57]), rather than the CT model, should be used to interpolate the data to $\rho(S_{n})$. However, the parameterizations of the level density parameter $a$ and the energy back-shift $E_{1}$ from Refs. [57, 65], commonly adopted in analyses with the Oslo method, fail to reproduce the shape of the NLD across the wide range of slopes suggested by the Cd and Sn isotopes. This is an additional complication that does not allow us to reliably use these parameters to estimate the value of the spin-cutoff parameter at the neutron threshold, required for the normalization of the GSF with the Oslo method, as well as for the shape method. In the latter case, the $\sigma(S_{n})$ value enters the excitation-energy dependence of the spin-cutoff provided by Eq. (11) and can significantly alter the slope of the GSFs extracted with this method.
@@
Calculations were performed for a wide range of $\sigma(S_{n})$ values, from 3 to 7. For each value, the GSFs corresponding to transitions to the ground state ($D1$) and the cluster of the higher-lying states ($D2$) were extracted and scaled to the available photo-dissociation data, ($\gamma,xn$), for 115In at 9.5-10.5 MeV (see also Sec.III). Further, the slope of the Oslo method GSF and the NLD was adjusted to the shape method result, and the optimal parameters $a$ and $E_{1}$ reproducing the NLD with the BSFG formula were estimated. The scaling factor, adjusting the Oslo method strength to the shape method result, estimated for each $\sigma(S_{n})$ is shown in Fig. 3. The range of the reduction factors corresponding to the 95% confidence interval for $\langle\Gamma_{\gamma}\rangle$ extracted from the systematics is indicated in Fig. 3 by the horizontal shaded area. This range was used to determine the range of $\sigma(S_{n})$ values and the recommended value of $\sigma(S_{n})$ used to extract the final results (vertical shaded area in Fig. 3). The recommended value corresponds to a reduction factor of 1.0, meaning that the Oslo and the shape method strengths align in absolute values after the slope adjustment, without any additional scaling. The recommended values of $\rho(S_{n})$, $D_{0}$, $a$, and $E_{1}$, along with their uncertainties, were estimated for the recommended value of $\sigma(S_{n})$ and its uncertainty range. All parameters determined in this procedure, as well as other parameters required for the normalization of the NLD are shown in Table 1.
@@ III.1 Nuclear level density of 109In
Figure 6: Experimental NLD of 109In, plotted together with a fit with the composite formula [56], CT, BSFG, Generalized Superfluid (GSM) phenomenological models, as well as calculations within the Skyrme-HF-BCS plus statistical [67], Skyrme-HFB plus combinatorial [53], and Gogny-HFB plus combinatorial [54] approaches (see Ref. [68] for more details regarding model calculations).
@@
A comparison of the experimental NLD with phenomenological and microscopic models, commonly adopted for calculations with the reaction code talys (version 2.0) [68], is shown in Fig. 6. All models, except for the Skyrme-HF-BCS plus statistical calculations [66], suggest much lower absolute values of the NLD below the neutron separation energy. The Skyrme-HF-BCS plus statistical model agrees quite well with the experiment up to approximately 6 MeV and deviates from it at higher energies. Analogous results are observed for the lightest studied 105Cd isotope (not shown in the figure). In general, very little experimental information is currently available on NLDs of unstable nuclei, except for a few cases studied with the Oslo and the $\beta$-Oslo methods [71]. In this respect, the 109In data presented in this work provide new valuable input for assessing the performance of different model approaches in the region of neutron-deficient heavy nuclei.
@@ III.2 $\gamma$-ray strength function of 109In
The GSF of 109In extracted with the Oslo method and the shape method is shown in Fig. 7. The GSF obtained with the shape method for $\gamma$-rays feeding the ground-state diagonal [D1 in Fig. 1(c)] and the cluster of four low-lying states around 1.1 MeV [D2 in Fig. 1(c)] using the recommended value of the spin-cutoff from Table 1 is presented by red and blue bands in Fig. 7. The uncertainty of the recommended $\sigma(S_{n})$ value was included in the shown uncertainty bands. The strengths overlap perfectly down to 4 MeV, chosen as the lower excitation-energy limit for the extraction of the GSFs. The shape method results cover a wide range of $\gamma$-ray energies (from 4 MeV to 10.4 MeV) and provide stringent constrains on the slope of the Oslo method GSF, thus yielding relatively small systematic uncertainties. The orange bands in Fig. 7 correspond to the total uncertainties, including statistical uncertainties. After the slope adjustment, the overall shape of the Oslo method strength is in excellent agreement with the shape method GSF, in line with the adopted Brink-Axel hypothesis. A slight deviation, still within the uncertainty bands, of the GSF for the diagonal D2 from the Oslo method strength at approximately 3 MeV may be potentially attributed to transitions of $E2$ type, not accounted for in the shape method.
[…]
Figure 8: Experimental GSF of 109In, shown together with the dipole strength model predictions with the $E1$ component based on the generalized (GLO) and simple (SLO) Lorentzians [68], the hybrid model [16], Skyrme-HF-BCS-QRPA [75], Skyrme-HFB-QRPA [76, 77], temperature-dependent Skyrme-HFB-QRPA [76], temperature-dependent relativistic mean field (RMF)-QRPA [78], Gogny-HFB-QRPA [79], Simplified Modified Lorentzian (SMLO) [80]. (b) Same as in (a) for $\gamma$-ray energies between 1 MeV and 10 MeV.
@@
The strength of 109In agrees well with the  [^1] Cd and  [^1] Sn data from approximately 4 to 7 MeV. This generally good agreement with the strengths of the neighboring nuclei indirectly supports the normalization of the 109In data, in particular, the resulting slopes of the NLD and the GSF. At lower energies, while still agreeing with the  [^1] Sn data within their rather broad uncertainty band, 109In falls well in between  [^1] Cd and  [^1] Sn, exhibiting no signs of a low-energy enhancement, the so-called upbend, observed in the lighter Cd isotopes [22].
An interesting feature observed in the new 109In data is the lack of any apparent enhancement in the vicinity of 8 MeV, which is clearly seen in the Sn isotopes [72, 21] and hinted at in the Cd and Pd data [22, 23]. This feature is essentially independent of the normalization procedure and makes the GSF of 109In stand out, in terms of shape, from the neighboring Pd, Cd, Sn, and Sb isotopes. Due to the absence of this 8-MeV enhancement, associated with the PDR in the earlier works  [22, 23], some of the theoretical model calculations employed in talys are much more compatible with the experimental data for 109In than in other cases in this mass region (see, e. g., [28]). The comparison of phenomenological models and calculations within the quasi-particle random-phase approximation (QRPA) with the experiment is shown in Fig. 8. Here, all options listed in the figure legend, except for the Gogny-HFB-QRPA [79] and the most recent Skyrme-HFB-QRPA [77], were combined with the default talys $M1$ model (SMLO from Ref. [80]) to produce the total dipole strength for the comparison with the experimental data. The Gogny-HFB-QRPA and Skyrme-HFB-QRPA $E1$ strengths were combined with the corresponding $M1$ strengths obtained within the same frameworks. In all cases, an upbend prescribed in Ref. [68] was added to the shown calculations. The Gogny-HFB-QRPA and Skyrme-HF-BCS-QRPA approaches agree nicely with the Oslo data below the neutron threshold within the experimental uncertainties. Substantial deviation of other theoretical strengths from each other and the experimental results is still observed, highlighting the importance of our new data in constraining theoretical uncertainties related to the GSF models, in much the same way as for the NLDs.

### v2

@@ II.1 Experimental details and the Oslo method
Figure 1: (a) Experimental $p-\gamma$ coincidence data, (b) unfolded, and (c) primary matrices for 109In. Yellow dashed lines mark the neutron separation energy. The ranges of excitation and $\gamma$-ray energies used for the analysis with the Oslo method are shown with solid blue lines in (c). The gates on the ground state with spin-parity $9/2^{+}$ (D1) and a cluster of states with spin-parities $3/2^{-}$, $11/2^{+}$, $5/2^{+}$, $1/2^{+}$ (D2) used for the shape method are shown with cyan dashed lines in (c). Excitation-energy and $\gamma$-ray-energy bins are 124-keV wide.
@@
The extracted primary $\gamma$-ray distributions $P(E_{\gamma},E_{i})$ are further decomposed in terms of the $\gamma$-transmission coefficient $\mathcal{T}_{i\rightarrow f}$ (proportional to the GSF) and the NLD $\rho_{f}$ as
@@
As outlined in Ref. [39], the decomposition provides the functional forms of the solutions for the NLD and the $\gamma$-transmission coefficient defined in the general form as:
@@
With one set of solutions found, $\rho(E_{i}-E_{\gamma})$ and $\mathcal{T}(E_{\gamma})$, an infinite set of equivalent solutions $\tilde{\rho}(E_{i}-E_{\gamma})$ and $\tilde{\mathcal{T}}(E_{\gamma})$ can be constructed by applying the scaling factors $A$ and $B$ and modifying the slope of the resulting functions with the parameter $\alpha$. To single out the unique physical solutions for the NLD and the $\gamma$-transmission coefficient, these parameters must be constrained using additional experimental data.
II.2 Standard Oslo method procedures for the normalization of the NLD and GSF
For the majority of Oslo method analyses applied to stable nuclei, the parameters $A$, $B$, and $\alpha$ are determined using information on low-lying discrete levels, available in the ENSDF compilation [48], together with $s$- or $p$-wave neutron resonance data listed in the Reference Input Parameter Library [49] and the Atlas of Neutron Resonances [50]. This part of the analysis is described in great detail in earlier publications, e. g., in [25, 28, 21]. In this work, the normalization of the 109In data uses this standard procedure as a starting point, and we therefore briefly summarize it in this section. The additional procedures employed in this work to extract the NLD and the GSF of 109In are provided in the following section.
@@ II.2 Standard Oslo method procedures for the normalization of the NLD and GSF
These procedures allow us to determine the scaling $A$ and the slope $\alpha$ of the NLD, where the latter also gives the slope of the GSF. As a final step, the scaling $B$ of the GSF can be extracted using the average total radiative width $\langle\Gamma_{\gamma}\rangle$ from $s$- and/or $p$-wave neutron capture experiments [50, 49]. The expression for the average total radiative width $\langle\Gamma_{\gamma}(E_{x},J,\pi)\rangle$ from Ref. [62] [Eq. (3.1)] can be adapted to account for the states accessed through the capture of $\ell=0$ or $\ell=1$ neutrons. For example, for the $s$-wave case and a target nucleus with $J_{t}\neq 0$, the GSF $f(E_{\gamma})$ can be expressed through the value of $\langle\Gamma_{\gamma}\rangle$ as [25]:
$\langle\Gamma_{\gamma}\rangle=$ $\langle\Gamma(S_{n},J_{t}\pm 1/2,\pi_{t})\rangle=\frac{1}{2\rho(S_{n},J_{t}\pm 1/2,\pi_{t})}$
$\times\int_{E_{\gamma}=0}^{S_{n}}dE_{\gamma}E_{\gamma}^{3}f(E_{\gamma})\rho(S_{n}-E_{\gamma})$
@@
Here, we used the proportionality of the $\gamma$-transmission coefficient $\mathcal{T}(E_{\gamma})$ and the GSF $f(E_{\gamma})$, $\mathcal{T}(E_{\gamma})=2\pi E_{\gamma}^{3}f(E_{\gamma})$, for dipole transitions expected to dominate in the studied energy region [62]. Within the Oslo method, it is not possible to distinguish between magnetic ($M$) and electric ($E$) types of transitions, unless external experimental information on the $M1$ distribution is available. The GSF extracted with Eq. (9) is therefore the sum of $E1$ and $M1$ dipole strengths. The GSF is extrapolated to low (and in some cases high) $E_{\gamma}$ energies to calculate the integral in Eq. (9) as discussed in Ref. [25].
@@
II.3 Normalization of the NLD and GSF in 109In
With neutron resonance data at hand, the normalization of both the NLD and the GSF becomes a relatively straightforward process. However, they are not available for some stable and unstable nuclei close to the valley of stability, including 109In, due to the lack of suitable targets for neutron-capture experiments. To tackle the normalization of the 109In data, some additional steps, outlined in the following subsections, are required.
II.3.1 Extracting the normalization coefficients for 109In from the systematics
In some cases, systematics for neighboring nuclei can be used to constrain the values of $\rho(S_{n})$ and $\langle\Gamma_{\gamma}\rangle$ [22, 21]. For odd-$Z$ nuclei, however, this approach is somewhat challenging, as only a limited number of stable isotopes are available to establish reliable systematics. This problem was, for example, encountered in the recent analysis of 127Sb [28], which relied largely on neutron resonance data from the Sn and Te isotopic chains. Likewise, very little information is available on In isotopes (except for the stable 113In and the long-lived 115In), that could potentially be used in the normalization procedure for the neutron-deficient 109In.
For the In isotopes, located one proton below the $Z=50$ closure, an approach similar to that adopted for 127Sb is suggested by the abundant neutron resonance data available for Sn and Cd nuclei. The $\langle\Gamma_{\gamma}\rangle$ value for the odd-even 109In was therefore extracted from s-wave $\langle\Gamma_{\gamma}\rangle$ values for the even-odd Sn and Cd isotopes from Ref. [50]. In the present case, a $\chi^{2}$ fit with a quadratic function was applied to the data for even-odd Sn and Cd nuclei plotted as a function of mass number (see Fig. 2). The quadratic fit was chosen as it provides a significantly better fit to the experimental data than a linear function (in contrast to the case in Ref. [28]), while maintaining a minimal set of assumptions made regarding the observed trend. All values are shown together with the quadratic fit in Fig. 2.
@@ II.3.1 Extracting the normalization coefficients for 109In from the systematics
II.3.2 Constraining the GSF slope with the shape method
To reduce the uncertainty in the slope shared by the NLD and the GSF, the so-called shape method was proposed in Ref. [63]. This technique has been successfully applied in different mass regions [63, 58, 45, 64] to reconstruct the general shape of the GSF (and, therefore, the slope of the NLD) from data on the direct decay to low-lying states. The number of counts $N$ associated with these decays is confined within a diagonal ($D$) in the primary matrix. The number of counts for a given $E_{i}$ bin can be written in a general form as [63]:
@@ II.3.2 Constraining the GSF slope with the shape method
where $f(E_{\gamma})E_{\gamma}^{3}$ is proportional to the $\gamma$-transmission coefficient, $\sigma(E_{i},J_{i})$ is the population cross-section for a state with spin $J_{i}$ at excitation energy $E_{i}$, and $g(E_{i},J_{i})$ is the intrinsic spin distribution given by Eq. (5). In cases when a diagonal contains more than one state, the first sum runs over all spins $J_{f}$ of the final states. As in the Oslo method, parity equipartition is adopted in Eq. (12), and only dipole transitions are taken into account. This restricts the applicability of the shape method to excitation energies where dipole transitions are dominant. As a next step, the population cross section is assumed to be independent of spin $J_{i}$ in the chosen region of excitation energies $E_{i}$. For dipole decays from a set of initial states to the final states with spins $J_{f}$, confined within the diagonal $D$ at the excitation-energy bin $E_{i}$ in the primary matrix, the GSF can therefore be expressed as:
@@
II.3.3 Consolidating the shape method and the systematics for the normalization of 109In
To proceed with the extraction of the GSF with the shape method and to use it for the normalization of the Oslo method strength, the spin-cutoff parameter at the neutron separation energy is required. The $\sigma(S_{n})$ value enters the excitation-energy dependence of the spin-cutoff provided by Eq. (11) and can significantly alter the slope of the resulting GSFs.
As mentioned in Sec. II.2, Eq.(7) or Eq.(8) are usually adopted in the Oslo method analysis together with Eq.(11) to model the spin-cutoff parameter dependence on the excitation energy. In the present case, however, the parameterizations of the level density parameter $a$ and the energy back-shift $E_{1}$ from Refs. [57, 65], commonly used in Eq.(7) and Eq.(8), cannot be reliably applied. The functional form of the NLD extracted from the decomposition in Eq. (1) exhibits a clear BSFG trend, however, the $a$ and $E_{1}$ parameters from Refs. [57, 65] fail to reproduce the shape of the NLD across the wide range of slopes suggested by the Cd and Sn isotopes. This is an additional complication that does not allow us to use these parameters to estimate the $\sigma(S_{n})$ value in a well-justified manner.
@@ II.3.3 Consolidating the shape method and the systematics for the normalization of 109In
Calculations were performed for a wide range of $\sigma(S_{n})$ values, from 3 to 7. For each value, the GSFs corresponding to transitions to the ground state ($D1$) and the cluster of the higher-lying states ($D2$) were extracted and scaled to the available photo-dissociation data, ($\gamma,xn$), for 115In at 9.5-10.5 MeV (see also Sec.III). Further, the slope of the Oslo method GSF and the NLD was adjusted to the shape method result through a $\chi^{2}$ minimization, and the optimal BSFG parameters $a$ and $E_{1}$ reproducing the NLD with this slope were estimated. The scaling factor, adjusting the Oslo method strength to the shape method result, estimated for each $\sigma(S_{n})$ is shown in Fig. 3. The range of these factors corresponding to the 95% confidence interval for $\langle\Gamma_{\gamma}\rangle$ extracted from the systematics is indicated in Fig. 3 by the horizontal shaded area. This range was used to determine the range of $\sigma(S_{n})$ values and the recommended value of $\sigma(S_{n})$ used to extract the final results (vertical shaded area in Fig. 3). The recommended value corresponds to a reduction factor of 1.0, meaning that the Oslo and the shape method strengths align in absolute values after the slope adjustment, without any additional scaling. The recommended values of $\rho(S_{n})$, $D_{0}$, $a$, and $E_{1}$ were also estimated for the recommended value of $\sigma(S_{n})$. Their uncertainty ranges are directly determined by the span of $\sigma(S_{n})$ values corresponding to the chosen $\langle\Gamma_{\gamma}\rangle$ scaling factor range (see Fig. 3). All parameters determined in this procedure, as well as other parameters required for the normalization of the NLD are shown in Table 1.
@@ III.1 Nuclear level density of 109In
Figure 6: (a) Experimental NLD of 109In plotted together with a fit with the composite formula [56], CT, BSFG, Generalized Superfluid (GSM) phenomenological models, as well as calculations within the Skyrme-HF-BCS plus statistical [66], Skyrme-HFB plus combinatorial [53], and Gogny-HFB plus combinatorial [54] approaches (see Ref. [67] for more details regarding model calculations). No adjustments were made to the latter three model calculations. (b) Experimental NLD of 109In plotted together with the adjusted Skyrme-HF-BCS, Skyrme-HFB, and Gogny-HFB models (see text.)
@@
A comparison of the experimental NLD with phenomenological and microscopic models, commonly adopted for calculations with the reaction code talys (version 2.0) [67], is shown in Fig. 6(a). No additional adjustments were applied to the shown microscopic calculations. All of them, except for the Skyrme-HF-BCS plus statistical calculations [66], suggest much lower absolute values of the NLD below the neutron separation energy. The Skyrme-HF-BCS plus statistical model agrees quite well with the experiment up to approximately 6 MeV and deviates from it at higher energies. Analogous results are observed for the lightest studied 105Cd isotope (not shown in the figure).
The three microscopic NLDs shown in Fig. 6(a) can be further adjusted to achieve better agreement with the experimental data. This was done by modifying the NLD slope and adding an energy shift according to Eq.(230) from Ref. [67]:
$\rho_{\rm mod}(E_{x},J,\pi)=\exp(c\sqrt{E_{x}-\delta})\rho(E_{x}-\delta,J,\pi).$ (15)
Here, $\rho_{\rm mod}$ is the modified NLD, $c$ and $\delta$ are the modifying parameters. By requiring the best fit to the experimental data from 2.5 MeV to $S_{n}(^{109}$In)=10.439=10.439 MeV, the following coefficients were found: $c=0.33(5)$ MeV-1/2, $\delta=0.37(4)$ MeV for the Skyrme-HF-BCS model, $c=0.64(4)$ MeV-1/2, $\delta=0.36(4)$ MeV for the Skyrme-HFB model, $c=0.54(4)$ MeV-1/2, $\delta=-0.54(2)$ MeV for the Gogny-HFB model. The adjustment of the latter model results in the best reproduction of the experimental data compared to all other considered theoretical approaches. In general, the modified Skyrme-HF-BCS and Gogny-HFB models reproduce the experimental data well from 2.5 MeV up to $S_{n}$ and support the $\rho(S_{n})$ value found in the normalization procedure using the shape method. All adjusted calculations are shown Fig. 6(b) together with the experimental NLD.
In general, very little experimental information is currently available on NLDs of unstable nuclei, except for a few cases studied with the Oslo and the $\beta$-Oslo methods [70]. In this respect, the 109In data presented in this work provide new valuable input for assessing the performance of different model approaches in the region of neutron-deficient heavy nuclei.
@@ III.2 $\gamma$-ray strength function of 109In
The GSF of 109In extracted with the Oslo method and the shape method is shown in Fig. 7. The GSF obtained with the shape method for $\gamma$-rays feeding the ground-state diagonal [D1 in Fig. 1(c)] and the cluster of four low-lying states around 1.1 MeV [D2 in Fig. 1(c)] using the recommended value of the spin-cutoff from Table 1 is presented by red and blue bands in Fig. 7. The uncertainty of the recommended $\sigma(S_{n})$ value was included in the shown uncertainty bands. The strengths overlap perfectly down to 4 MeV, chosen as the lower excitation-energy limit for the extraction of the GSFs. The shape method results cover a wide range of $\gamma$-ray energies (from 4 MeV to 10.4 MeV) and provide stringent constrains on the slope of the Oslo method GSF, thus yielding relatively small systematic uncertainties. The orange bands in Fig. 7 correspond to the total uncertainties, including statistical uncertainties. After the slope adjustment, the overall shape of the Oslo method strength is in excellent agreement with the shape method GSF, in line with the adopted Brink-Axel hypothesis. A slight deviation, still within the uncertainty bands, of the GSF for the diagonal D2 from the Oslo method strength at approximately 3 MeV may potentially be attributed to transitions of $E2$ type, not accounted for in the shape method.
Figure 7: (a) Oslo method GSF of 109In (orange band), shown together with the shape method results (red and blue shaded bands), the strength from the Coulomb excitation ($p,p^{\prime})$ experiment on 112Sn [71], and photo-dissociation ($\gamma,xn$) data for 115In [72, 73]. (b) Same as in (a) for $\gamma$-ray energies between 1 MeV and 10 MeV (the Oslo method data points are shown).
Figure 8: (a) Comparison of the Oslo method GSF of 109In (orange band with squared markers) with the Oslo method strengths of the neighboring Sn isotopes,  [^1] Sn and 112Sn from the ($p,d\gamma$) [31, 21] and ($p,p^{\prime}\gamma$) [74] experiments, respectively, shown together with the Coulomb excitation data for 112Sn [71] and photo-dissociation ($\gamma,xn$) data for 115In [72, 73]. (b) Comparison of the Oslo method 109In result with the 106, [^1] Cd [22] and 108Pd [23] Oslo method strengths from the 3He-induced reactions, as well as the ($\gamma,xn$) data for 115In.
Figure 9: Experimental GSF of 109In extracted with the Oslo and shape methods, shown together with the dipole strength model predictions with the $E1$ component based on the generalized (GLO) and simple (SLO) Lorentzians [67], the hybrid model [16], Skyrme-HF-BCS-QRPA [75], Skyrme-HFB-QRPA [76, 77], temperature-dependent Skyrme-HFB-QRPA [76], temperature-dependent relativistic mean field (RMF)-QRPA [78], Gogny-HFB-QRPA [79], Simple Modified Lorentzian (SMLO) [80]. (b) Same as in (a) for $\gamma$-ray energies between 1 MeV and 10 MeV.
@@
The comparison with other experimental data, in particular the earlier Oslo method results, is shown in Fig. 8. Here, the shape method strengths for 109In are omitted as they are in excellent agreement with the Oslo method data. The 109In and the  [^1] Sn [31, 21] strengths are in fairly good agreement from approximately 4 to 7 MeV within the uncertainty bands. The recent result by Söderström et al. [74] exhibits a similar shape but a somewhat higher strength across the whole shown energy range. An overall good agreement above 5 MeV is also observed between 109In and the closest earlier studied 106, [^1] Cd isotopes [22]. The Pd data from the study by Eriksen et al. [23] are also shown in Fig. 8(b). The shapes of GSFs of the 105-108Pd isotopes from this work are quite similar, and, therefore, only the 108Pd case is shown here. These data agree quite well with 109In at low energies but, similar to the new 112Sn result [see Fig. 8(a)], demonstrate a slightly larger strength at higher energies.
This generally good agreement with the strengths of the neighboring nuclei indirectly supports the normalization of the 109In data, in particular the resulting slopes of the NLD and the GSF. At lower energies, while still agreeing with the  [^1] Sn data within their rather broad uncertainty band, 109In falls well in between the Cd and Sn cases, exhibiting no signs of a low-energy enhancement, the so-called upbend, observed in the lighter Cd isotopes [22].
An interesting feature observed in the new 109In data is the lack of any apparent enhancement in the vicinity of 8 MeV, which is clearly seen in the Sn isotopes (see Fig. 7, Fig. 8(a), and Refs. [71, 21]) and hinted at in the Cd and Pd data (see Fig. 8(b) and Refs. [22, 23]). This feature is essentially independent of the normalization procedure and makes the GSF of 109In stand out, in terms of shape, from the neighboring Pd, Cd, Sn, and Sb isotopes. Due to the absence of this 8-MeV enhancement, associated with the PDR in the earlier works  [22, 23], some of the theoretical model calculations employed in talys are much more compatible with the experimental data for 109In than in other cases in this mass region (see, e. g., [28]). The comparison of phenomenological models and calculations within the quasi-particle random-phase approximation (QRPA) with the experiment is shown in Fig. 9. Here, all options listed in the figure legend, except for the Gogny-HFB-QRPA [79] and the most recent Skyrme-HFB-QRPA [77], were combined with the default talys $M1$ model (SMLO from Ref. [80]) to produce the total dipole strength for the comparison with the experimental data. The Gogny-HFB-QRPA and Skyrme-HFB-QRPA $E1$ strengths were combined with the corresponding $M1$ strengths obtained within the same frameworks. In all cases, an upbend prescribed in Ref. [67] was added to the shown calculations. The Gogny-HFB-QRPA and Skyrme-HF-BCS-QRPA approaches agree nicely with the Oslo data below the neutron threshold within the experimental uncertainties. Substantial deviation of other theoretical strengths from each other and the experimental results is still observed, highlighting the importance of our new data in constraining theoretical uncertainties related to the GSF models, in much the same way as for the NLDs.